A flat piece of ice floats in calm seawater (density = 1025 kg/m3) and protrudes 3.25 cm above the surface. What is the minimum surface area so that an 84.9-kg man can stand on this ice without his feet getting wet?
1
Expert's answer
2018-06-21T10:19:08-0400
Mg=(ρ_seawater-ρ_ice )gV
M=(ρ_seawater-ρ_ice )hA
A=M/(ρ_seawater-ρ_ice )h=84.9/(1025-917)0.0325=24.2 m^2.